OpenStudy (anonymous):

Determine the interior, the boundary and the closure of the set {z ε: Re(z^2>1} Is the interior of the set path-connected?

OpenStudy (anonymous):

Alright so z^2=(x+iy)(x+iy)=x^2+2ixy-y^2 so Re(x^2+2ixy-y^2)= x^2-y^2 >1 So would the image be a hyperbola that starts when each axis is >1 leaving a hole in the center? Boundary: {z ε: Re(z^2)=1} Interior: none Closure: {z ε: Re(z^2)>1} Since there is no interior the question of interior path connectedness is mout. I'm not sure if this is right though @eliassaab @experimentX @UnkleRhaukus @phi @eseidl

OpenStudy (ybarrap):

Wouldn't the interior be R(^2) > 1 because for there exist an r > 0 such that $$ B_r(x)\subset R(z^2)> 1 $$ for all x in R(z^2) > 1?

OpenStudy (ybarrap):

Boundary looks right.

OpenStudy (anonymous):

That helps! Thanks!

OpenStudy (ybarrap):

you're welcome